Vector P has magnitude of 6 units in the direction of positive x-axis. Vector Q has a magnitude of 6√3 Units and lies in x-y plane, making an angle of 30° with positive x-axis and 60° with positive y-axis. Find the vector product of PxQ

1
Expert's answer
2021-12-15T16:22:07-0500

"\vec P \times\vec Q=|\vec P|\cdot|\vec Q|\cdot\sin 30^\circ=6\cdot6\sqrt{3}\cdot\sin 30^\circ=31.2\ (units)"